# Atomistic Framework for Glassy Polymer Viscoelasticity Across 20 Frequency Decades

### v1

@@ Abstract
In Fig. 3, we plot the shear modulus of PMMA as a function of the external frequency $\Omega$ at $T=300$ K, representing room temperature conditions. The results of the NALD atomistic calculation is compared with MD results, our experimental DMA data as well as literature reported from different mechanical testing techniques, across a broad frequency range. Four distinct frequency regimes, which correspond to structural processes within PMMA, are observed and reliably captured by NALD. First, at the high-frequency end (Petahertz region), a plateau emerges corresponding to the affine modulus which is a purely elastic contribution determined by the instantaneous bonding polymer chain or network; the second regime is the resonance frequency region (THz), where two sharp peaks appear: the first is at 300300  THz, and the second one is close to $600$ THz, due to the vibrational resonances of PMMA chemical bonds ali_2015review; the atomistic MD simulations accurately reproduce the resonance behavior. The third regime is the GHz regime, which coincides with experimental Brillouin spectroscopy results weishaupt1995pressure. Here, non-affine deformations become apparent in the shear modulus i.e. the first decay after the resonance peak indicates that the molecules begin to relax MULLERPABEL_2022_107701; joaquin_2025aging.
@@
At kHz frequencies, we include Split-Hopkinson test data MULLIKEN20061331; Sahraoui_1994; moy_paul_2003, noting that the strain rate used to define the testing speed is not directly equivalent to a frequency; however, the stress wave generated during testing typically lies in the kHz range. Some experimental data agree well with NALD predictions, whereas others overestimate/underestimate the modulus. Since Split-Hopkinson pressure bars are custom-built and testing conditions vary, such deviations are not unexpected. At lower frequencies, available techniques are limited to contact-based methods, each with its own limitations, leading to additional scatter. Literature shows that, even for a well-studied glassy polymer such as PMMA, an exact modulus is not possible to obtain from classical tests alone christofl2021comprehensive. Measurements spanning nano- to macroscales demonstrate that local material variations and experimental constraints (e.g., deformation mode, sample preparation) only allow one to define a range of values, with nanoscale AFM experiments typically yielding lower moduli christofl2021comprehensive; ganser2018combining. In our experimental study, we use PMMA from the same manufacturer for DMA experiments in the low-frequency regime (0.1–100 Hz). The DMA results match the NALD estimates and reveal a secondary decay in the shear modulus associated with the $\beta$-relaxation process MULLIKEN20061331; moy_paul_2003; Sahraoui_1994.
@@
To summarize, we provided the atomic-scale computation of the frequency-dependent shear modulus of a glassy polymer (PMMA) across a broad frequency spectrum, from several hundreds of terahertz down to millihertz. The computation scheme allows for multiscale bridging in time, thanks to the calculation of non-affine atomic displacements, which contribute substantial negative corrections to the modulus, as the deformation frequency is lowered. The calculations are successfully compared with MD simulations of the deformation process (limited to the terahertz regime), as well as with several sets of experimental data comprising both Brillouin light scattering (at GHz) and mechanical testing (from $10^{5}$ Hz down to millihertz). While a parameter-free quantitative match is possible, in order to capture the secondary $\beta$ relaxation in the modulus that occurs at hundreds of Hertz, a power-law memory function for the atomic-scale friction kernel is required, in agreement with previous results in the literature milster_2024tracer; li2017computing. Overall, the proposed NALD framework allows us to bridge the timescale gap between non-destructive light scattering techniques and conventional mechanical testing of real soft materials, and provides a theoretical, experimentally validated description of the mechanics of glassy polymers across 20 orders of magnitude in deformation frequency, with profound implications for both fundamental science and engineering applications. A future challenge will be to extend this atomistic framework to the vicinity of the glass transition glova.
@@
Acknowledgments- AS, VV, and AZ gratefully acknowledge funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. VV acknowledges the computational resource provided via the project “TPLAMECH” on INDACO platform at the HPC facility of Università degli Studi di Milano. AZ gratefully acknowledges funding from the US Army DEVCOM Army Research Office through contract nr. W911NF-22-2-0256. CC acknowledges the Hertha Firnberg program (project no. T 1314-N) of the Austrian Science Fund (FWF), Grant DOI: 10.55776/T1314 for funding. Discussions with Simone Napolitano and Federico Caporaletti are gratefully acknowledged.

### v2

@@ Abstract
In Fig. 3, we plot the shear modulus of PMMA as a function of the external frequency $\Omega$ at $T=300$ K, representing room temperature conditions. The results of the NALD atomistic calculation are compared with MD results, our experimental DMA data (details in section S3 of SI 1) as well as literature data from different mechanical testing techniques, across a broad frequency range. Four distinct frequency regimes, which correspond to structural processes within PMMA, are observed and reliably captured by NALD. First, at the high-frequency end (above terahertz region), a plateau emerges corresponding to the affine modulus, a purely elastic contribution determined by the instantaneous bonding polymer chain or network; the second regime is the resonance frequency region (THz), where two sharp peaks appear: the first is at 300300  THz, and the second one is close to $600$ THz, due to the vibrational resonances of PMMA chemical bonds Ali et al. (2015); the atomistic MD simulations accurately reproduce the resonance behavior. The third regime is the GHz regime, which coincides with experimental Brillouin spectroscopy results Weishaupt et al. (1995). Here, non-affine deformations become significant. The first decay in modulus arises from non-affine relaxations that reduce the affine contribution to the shear modulus, resulting in mechanical softening Müller-Pabel et al. (2022); Joaquín-Jaime et al. (2025).
@@
At kHz frequencies, we include Split-Hopkinson test data Mulliken and Boyce (2006); Sahraoui and Lataillade (1994); Moy et al. (2003), noting that the strain rate used to define the testing speed is not directly equivalent to a frequency; however, the stress wave generated during testing typically lies in the kHz range. Some experimental data agree well with NALD predictions, whereas others either overestimate or underestimate the modulus (For strain-rate and modulus conversions, see Section S4 of the SI 1). Since Split-Hopkinson pressure bars are custom-built and testing conditions vary, such deviations are not unexpected. At lower frequencies, available techniques are limited to contact-based methods, each with its own limitations, leading to additional scatter. Literature shows that, even for a well-studied glassy polymer such as PMMA, an exact modulus is not possible to obtain from classical tests alone Christöfl et al. (2021). Measurements spanning nano- to macroscales demonstrate that local material variations and experimental constraints (e.g., deformation mode, sample preparation) only allow one to define a range of values, with nanoscale AFM experiments typically yielding lower moduli Christöfl et al. (2021); Ganser et al. (2018). In our experimental study, we use PMMA from the same manufacturer for DMA experiments in the low-frequency regime (0.1–100 Hz) (details in section S3 of SI 1). The DMA results match the NALD estimates and reveal a secondary decay in the shear modulus associated with the $\beta$-relaxation process Mulliken and Boyce (2006); Moy et al. (2003); Sahraoui and Lataillade (1994).
@@
To summarize, we provided the atomic-scale computation of the frequency-dependent shear modulus of a glassy polymer (PMMA) across a broad frequency spectrum, from several hundreds of terahertz down to millihertz. The computation scheme enables multiscale bridging in time, thanks to the calculation of non-affine atomic displacements, which contribute substantial negative corrections to the modulus, as the deformation frequency decreases. The calculations are successfully compared with MD simulations of the deformation process (limited to the terahertz regime), as well as with several sets of experimental data comprising both Brillouin light scattering (at GHz) and mechanical testing (from $10^{5}$ Hz down to millihertz). While a near-parameter-free quantitative match is possible, in order to capture the secondary $\beta$ relaxation in the modulus that occurs at hundreds of Hertz, a power-law memory function for the atomic-scale memory kernel is required, in agreement with previous results in the literature Milster et al. (2024); Li et al. (2017). Overall, the proposed NALD framework allows us to bridge the timescale gap between non-destructive light scattering techniques and conventional mechanical testing of real soft materials, and provides a theoretical, experimentally validated description of the mechanics of glassy polymers across twenty orders of magnitude in deformation frequency, with profound implications for both fundamental science and engineering applications. A future challenge will be to extend this atomistic framework to the vicinity of the glass transition Glova and Karttunen (2024) and to more complex material architectures Varol et al. (2020).
Data Availability
The data that support the findings of this study are available within the article and its Supplementary Information.
Acknowledgements
@@
AS, VV, and AZ gratefully acknowledge funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. VV acknowledges the computational resource provided via the project “TPLAMECH” on INDACO platform at the HPC facility of Università degli Studi di Milano. AZ gratefully acknowledges funding from the US Army DEVCOM Army Research Office through contract nr. W911NF-22-2-0256. CC acknowledges the Hertha Firnberg program (project no. T 1314-N) of the Austrian Science Fund (FWF), Grant DOI: 10.55776/T1314 for funding. Discussions with Simone Napolitano and Federico Caporaletti are gratefully acknowledged.